OpenStudy (anonymous):

the sum of two numbers is 12. find the minimum value of the sum of their squares. find the smaller number such that the product of one by the square of the other is to be a maximum . find the bigger number such that the product of one by the cube of the other is to be a maximum

OpenStudy (tkhunny):

a and b are the numbers So a + b = 12 Minimize a^2 + b^2 = a^2 + (12-a)^2 Go!